OpenStudy (anonymous):

50 is 6.25% of what number?

OpenStudy (anonymous):

divide 50 by 6.25 to get 1%, then multiply by 100 to get 100% (original)

mathslover (mathslover):

Let that number be x. So: According to the statement it is clear that : \(50 = 6.25 \% \times x \) \(50 = \cfrac{6.25}{100} \times x\)

mathslover (mathslover):

Solve for x now.

OpenStudy (anonymous):

16?

mathslover (mathslover):

No, @CleoXoXo98 , try again.

mathslover (mathslover):

CleoXoXo98 , may I know how you got 16?

OpenStudy (anonymous):

i did what i wasn't supposed to i just realized haha

mathslover (mathslover):

Haha! Try again cleoXoXo98 but this time, try what you were supposed to do.

OpenStudy (anonymous):

idk :(

OpenStudy (anonymous):

oh wait...8?

mathslover (mathslover):

No Cleo. Where are you having problem?

OpenStudy (anonymous):

i just suck at math that's my problem :(

mathslover (mathslover):

No problem. See : I have \(50 = \cfrac{6.25}{100} \times x\) right?

OpenStudy (anonymous):

mhm

mathslover (mathslover):

Ok! No multiply 100 both sides.

OpenStudy (anonymous):

k

mathslover (mathslover):

What do you get?

OpenStudy (anonymous):

625 but i feel like that's wrong

OpenStudy (anonymous):

i feel really dumb right about now :/